OpenStudy (rainbow_rocks03):

Will medal someone for the best response so help me plz. The area of a rectangle is 7 2/5 square feet. The length of this rectangle is 3 1/12 feet. What is the width?

OpenStudy (rainbow_rocks03):

PLZ HELP ME ASAP! someone

OpenStudy (bradely):

Area of the rectangle A = length *width 7 (2/5) =3(1/2) *width 37/5 = 7/2 *width width = (37/5)/(7/2) = (37/5)(2/7) =74/35 = 2 (4/35)feet Source: http://www.mathskey.com/question2answer/

OpenStudy (rainbow_rocks03):

SOMEONE PLZ HELP ME!

OpenStudy (rainbow_rocks03):

here are the choices. A.1 5/12 B.5/12 C.2 2/5 D.2 1/3

OpenStudy (bradely):

Area of the rectangle A = length *width 7 (2/5) =3(1/12) *width 37/5 = 37/2 *width width = (37/5)/(37/12) = (37/5)(12/37) =12/5 = 2 (2/5)feet option C is correct Source: http://www.mathskey.com/question2answer/

OpenStudy (rainbow_rocks03):

thank you SO MUCH!